OpenStudy (anonymous):

If y varies directly as x, and y = 15 as x = -3, find y for the x-value of 4.

OpenStudy (anonymous):

Direct variation takes the form y=kx. to find k, plug in the given values: 15 = k(-3) k = -5 now you have y = -5x, so plug in 4 for x to find y: y = -5(4) = -20 inverse variation takes the form y=k/x. find k: 5 = k/6 k = 30 formula: y = 30/x when x=10: y = 30/10 = 3